OpenStudy (anonymous):

solve the problem exactly: log(2x+7)=1+log(x-1

OpenStudy (dumbcow):

make both sides as power of 10 \[10^{\log_{}2x+7 }=10^{1+\log_{}x-1 }\] this will cancel the logs 2x+7 = 10(x-1) -8x = -17 x = 17/8

OpenStudy (anonymous):

thanks :)